# Evaluating Visual and Cultural Interpretation: The K-Viscuit Benchmark with Human-VLM Collaboration

### v1

@@ 4 Experiments
Model All Food Beverage Game Celeb. Religion Tool Clothes Heritage Arch. Agri. InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19
@@ 4.1 Models
The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), and Idefics2-8B Laurençon et al. (2024). We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), and Gemini-1.5-Pro (Reid et al., 2024). All models are evaluated in the conventional multiple-choice setup, where a model is prompted to choose its answer from one of four options. The input text is constructed by concatenating (1) a question, (2) each option with option letters in alphabetical order, and (3) the instruction about output format (i.e., "Answer with the option’s letter from the given choices directly."). We use accuracy as an evaluation metric.
@@
4.2 Results
Table 2 demonstrates the evaluation results of different VLMs on our dataset. We first observe that proprietary models usually show higher accuracies. The Gemini-1.5-Pro and GPT-4-Turbo achieve the highest and second-highest scores, respectively. Among the open-sourced models, Idefics2 and InternLM-XComposer2-VL usually perform better than other models. The accuracy of models in different question types is shown in Table 3. We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions, except for Gemini-1.5-Pro and Idefics2-8B. Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs. Our dataset enables the evaluation and identification of such aspects where models can be improved.
@@ 4.3 Analyses
Model en ko en + ko
Claude-3-Opus 70.02 65.30 70.32
GPT-4-Turbo 80.82 75.34 80.37
Gemini-1.5-Pro 81.58 80.82 83.41
@@
To evaluate how well people from different backgrounds perform on K-Viscuit, we conducted human evaluation. We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions. This test was administered to 20 Koreans and 14 Non-Koreans. As depicted in Fig. 6, the results showed that participants with better knowledge of Korean culture achieved higher accuracy. However, even among Koreans, knowledge gaps exist both between individuals and within a single person’s knowledge across different domains, such as history. Proprietary VLMs can match or exceed human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators. Details of the user study are provided in the Appendix B.2.
Figure 6: Human Evaluation. Accuracy comparison between Koreans and non-Koreans on a K-Viscuit subset, measured across different proficiency levels.
@@
Fig. 5 presents the prediction of selected models on sampled examples. In the first example, all models fail to correctly answer the question about asking detailed game rules. In the second case, two proprietary models are wrong while open-source models make correct answers. The third problem requires the model to recognize the structure in the image as Cheomseongdae and infer that it serves a similar function to the Griffith Observatory in the United States. Both proprietary models make correct answers to this example. In the final example, all models successfully identify the correct answer about agriculture.
@@ 5.1 Evaluation beyond Multiple-choice VQA
Our dataset is designed as a multiple-choice VQA task, where VLMs are required to choose one of the given options. While this classification setup is advantageous for measuring model performance through accuracy, it may not fully capture the models’ understanding of Korean culture. For instance, VLMs may correctly select an answer from the given options but fail in a generative VQA setup where they must generate a free-form answer. An example of this is shown in Fig. 7 (left), where the VLM correctly identifies an object from the choices but hallucinates it as “stones or bricks” when generating an answer without options. In contrast, the right example in Fig. 7 shows the VLM accurately identifying the object as “bibimbap” in both multiple-choice and generative setups. Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
@@ 5.2 Augmenting VLMs with External Knowledge
In previous experiments, we observed significant performance gaps in models for questions that require cultural knowledge. To address these gaps, two methodologies can be considered: fine-tuning open-source models and leveraging external knowledge. While fine-tuning is limited to open-source models, leveraging external knowledge can be applied to both open-source and proprietary models. Therefore, we focused on the latter approach.
We augmented the models with relevant documents for each test image from the Food concept. Our retrieval method involved generating captions using GPT-4-Turbo, embedding these captions via text-embedding-3-large to build queries, and retrieving Top-1 document from 152 Wikipedia pages related to the objects mentioned in the Type 1 options. The distractors in K-Viscuit were designed to be highly similar to the correct answers. This creates a challenging retrieval setting with numerous hard negatives, aiming to mimic real-world conditions. We aim to assess whether the retrieval method remains effective even in such a setting by observing the performance changes when cultural knowledge is added. The results show that providing retrieved documents can improve model performance, as indicated by higher scores in the Retrieved setting compared to None. The Oracle setting, which uses a curated relevant document, serves as an upper bound. Overall, augmenting VLMs with external knowledge via document retrieval demonstrates the potential to bridge the performance gap in VQA tasks requiring cultural context. We provide the details in the Appendix D.
Model None Retrieved Oracle
LLaVA-1.6-7B 43.66 68.31 78.87
LLaVA-1.6-13B 45.77 64.08 80.28
InternLM-XC2-7B 50.70 68.31 82.39
Idefics2-8B 51.41 67.61 82.39
Claude-3-opus 62.68 70.42 87.32
GPT-4-Turbo 73.94 78.17 88.73
Gemini-1.5-Pro 80.28 78.17 90.85
@@ 6 Conclusion
In this work, we proposed a semi-automated pipeline to construct culturally aware benchmarks for vision-language models, featuring multi-choice QA with two types of questions that require visual recognition or reasoning. This pipeline combines human-VLM collaboration, enhancing efficiency and diversity in generating culturally relevant QAs. Our approach, first used to create the K-Viscuit dataset, can be adapted to other cultures.
Our evaluation with K-Viscuit identified a notable performance gap between open-source and proprietary VLMs in understanding Korean culture, highlighting the need for improvement in open-source models. Further analyses demonstrated the impact of in-depth cultural understanding on VQA results. Finally, we enhanced VLM performance through knowledge augmentation and explored open-ended answer generation beyond the multiple-choice setting. Our work stresses the importance of cultural diversity in evaluation, aiming for more inclusive and accurate VLMs in the future.

### v2

@@ 4.1 Models
The following open-source VLMs are used for experiments: InstructBLIP-7B/13B (Dai et al., 2024), LLaVA-v1.6-7B/13B (Liu et al., 2024), mPLUG-Owl2-7B (Ye et al., 2023), InternLM-XComposer2-VL-7B (Dong et al., 2024), Molmo-7B-D (Deitke et al., 2024), Idefics2-8B Laurençon et al. (2024), and Llama-3.2-11B-Vision-Instruct (Dubey et al., 2024). We also use the following proprietary models: Claude-3-opus (Anthropic, 2024), GPT-4-Turbo (Achiam et al., 2023), Gemini-1.5-Pro (Reid et al., 2024), and GPT-4o (OpenAI, 2024). All models are evaluated in the conventional multiple-choice setup, where a model is prompted to choose its answer from one of four options. The input text is constructed by concatenating (1) a question, (2) each option with option letters in alphabetical order, and (3) the instruction about output format (i.e., "Answer with the option’s letter from the given choices directly."). We use accuracy as an evaluation metric.
4.2 Results
Table 2 demonstrates the evaluation results of different VLMs on our dataset. We first observe that proprietary models usually show higher accuracies. The GPT-4o and Gemini-1.5-Pro achieve the highest and second-highest scores, respectively. Among the open-sourced models, Llama-3.2-11B and Idefics2-8B usually perform better than other models. The accuracy of models in different question types is shown in Table 3. We observe that most models show higher accuracy in Type 2 questions compared to Type 1 questions. Regarding these trends, we suspect visual recognition with diverse cultural contexts poses inherent challenges for VLMs. Our dataset enables evaluating and identifying such aspects where models can be improved.
@@ 4.2 Results
Model en ko en + ko
Claude-3-Opus 70.02 65.30 70.32
GPT-4-Turbo 80.82 75.34 80.37
Gemini-1.5-Pro 81.58 80.82 83.41
GPT-4o 89.50 76.56 79.76
@@ 4.3 Analyses
Figure 5: Human Evaluation. Accuracy comparison between Koreans and non-Koreans on 50 samples of K-Viscuit. The performance of selected models on this subset is also displayed. The average scores for Koreans and non-Koreans are 80.2 and 47.0, respectively.
@@
We conducted a human evaluation to evaluate how well people from different backgrounds perform on K-Viscuit. We selected a subset of our dataset by randomly sampling 25 images, each with one type 1 and one type 2 question, totaling 50 questions. This test was administered to 20 Koreans and 14 non-Koreans. As depicted in Fig. 5, the results showed that participants with better knowledge of Korean culture achieved higher accuracy. However, even among Koreans, knowledge gaps exist between individuals and within a single person’s knowledge across different domains, such as history. A Proprietary VLM (i.e., GPT-4-Turbo) shows comparable performance to human performance, indicating that utilizing VLMs for generating diverse, culturally nuanced questions is more effective than relying solely on individual human annotators. Details of the user study are provided in the Appendix B.2.
VQAv2 CVQA K-Viscuit
LLaVA-v1.6-13B 82.8 57.9 57.1
Molmo-7B-D 85.6 65.5 61.0
Idefics2-8b 81.2 69.0 63.6
Llama-3.2-11B 75.2 72.4 68.0
Table 5: Evaluation results of selected open-source VLMs on various VQA datasets. The results on the VQAv2 (Goyal et al., 2017) are taken from the respective papers of each model. The performance on the CVQA dataset was measured on the Korean subset.
Comparison of open-source VLMs on Various VQA Datasets
We conducted experiments to measure the performance of VLMs on various VQA datasets. To this end, we compared the performance of four open-source VLMs on commonly used VQA benchmarks: the VQAv2 (Goyal et al., 2017) dev-test split, the Korean subset of CVQA (Romero et al., 2024), and our dataset, with results shown in Table 5. The VLMs demonstrated their best performance on VQAv2, while showing relatively lower accuracy on CVQA and our dataset. This suggests that the cultural questions we collected are indeed challenging for VLMs to solve.
@@
Fig. 6 presents the prediction of selected models on sampled examples. In the first example, all models fail to correctly answer the question about asking detailed game rules. In the second case, two proprietary models are wrong while open-source models make correct answers. The third problem requires the model to recognize the structure in the image as Cheomseongdae and infer that it serves a function similar to that of the Griffith Observatory in the United States. Both proprietary models make correct answers to this example. In the final example, all models successfully identify the correct answer about agriculture.
@@ 5.1 Evaluation beyond Multiple-choice VQA
Our dataset is designed as a multiple-choice Visual Question Answering (VQA) task, where Vision-Language Models (VLMs) select one answer from given options. While this classification setup enables straightforward performance measurement through accuracy, it may not fully reveal the models’ depth of cultural understanding. For instance, VLMs may correctly select an answer from the available options but fail in a generative VQA setup, where they must generate a free-form answer. As shown in Fig. 7, while the VLM accurately identifies the object as bibimbap in both multiple-choice and generative setups for the first example, it fails to provide accurate cultural context in the second example despite choosing the correct option in the multiple-choice format.
To quantitatively analyze this aspect, we randomly sampled 80 questions from the Food category and evaluated Llava-v1.6-13B’s performance in a generative setting. The model was prompted with the instruction: "This is a question about Korean culture. Answer the given questions briefly and concisely." without access to the original multiple-choice options. Two native Korean evaluators assessed the generated responses using three categories: (1) Correct: The model’s response fully contained the original answer or was reasonably accurate to the question, (2) Hallucinated: The model confidently provided incorrect information, and (3) Abstained: The model’s response did not include a direct answer to the question but provided a relevant description of the image and did not attempt to generate an answer hastily (e.g., "The specific type of food and its cultural context would determine when it is commonly served").
The human analysis results are shown in Table 6. We observed that the VLM’s accuracy in the generative setup was relatively lower than in the multiple-choice setup (from 45.07% to 36.25%). This decrease in accuracy is likely because the task is inherently more difficult without predefined options. Interestingly, a significant proportion of responses (21.25%) were classified as abstained, indicating that the VLM recognized the need for additional information to answer confidently and refrained from providing a definitive but potentially incorrect response. This behavior suggests that providing the necessary external knowledge could enhance performance. Our future research aims to propose various evaluation methods to assess whether VLMs genuinely possess the knowledge to answer questions accurately.
Correct Hallucinated Abstained Total 29 (36.35%) 34 (42.5%) 17 (21.25%) 80 (100%)
Table 6: Human evaluation results of LLaVA-1.6-13B on Food category.
@@ 5.2 Augmenting VLMs with External Knowledge
In previous experiments, models struggled with questions requiring cultural knowledge. To address these gaps, we considered fine-tuning open-source models and augmenting models with external knowledge. Since the latter can be applied to both open-source and proprietary models, we focused on that approach. We augmented the models with relevant documents for each test image from the Food concept. Our retrieval method involved generating captions using GPT-4-Turbo, embedding these captions via text-embedding-3-large to build queries, and retrieving Top-1 document from 152 Wikipedia pages related to the objects mentioned in the Type 1 options. The K-Viscuit distractors were designed to closely resemble correct answers, creating a challenging retrieval setting with numerous hard negatives that mimic real-world conditions. We assess if the retrieval method remains effective under these conditions by examining performance changes when cultural knowledge is introduced.
As shown in Table 7, Providing retrieved documents can enhance model performance, as demonstrated by improved scores in the Retrieved setting compared to None. In cases where proprietary models did not benefit from certain retrieved documents, their performance still surpassed the baseline when given carefully curated Oracle documents. This suggests that retrieval-based augmentation has the potential to strengthen cultural knowledge in VQA tasks, provided that the quality and relevance of the selected documents are refined. We provide the details in the Appendix D.
Model None Retrieved Oracle LLaVA-1.6-7B 43.66 68.31 78.87 LLaVA-1.6-13B 45.77 64.08 80.28 InternLM-XC2-7B 50.70 68.31 82.39 Idefics2-8B 51.41 67.61 82.39 Claude-3-opus 62.68 70.42 87.32 GPT-4-Turbo 73.94 78.17 88.73 Gemini-1.5-Pro 80.28 78.17 90.85 GPT-4o 88.73 83.10 92.25
@@ 6 Conclusion
In this work, we proposed a semi-automated framework for constructing culturally aware benchmarks for vision-language models through human-VLM collaboration, focusing on visual recognition and reasoning in multi-choice QA. We demonstrated its effectiveness by creating K-Viscuit, revealing a significant performance gap between proprietary and open-source VLMs in understanding Korean culture. Through detailed analyses and knowledge augmentation experiments, we established the impact of cultural understanding on VQA performance and extended our investigation to open-ended answer generation. This work highlights the importance of cultural diversity in model evaluation and paves the way for more inclusive VLMs.
@@
Ethical Statement
In constructing K-Viscuit, we ensured all images were sourced from Wikimedia Commons under Creative Commons licenses, maintaining proper attribution and copyright compliance. The dataset was carefully curated to avoid harmful or inappropriate content, and all questions and answers were verified by native Korean speakers to ensure cultural relevance. While comprehensive, we acknowledge that our dataset captures only a subset of Korean cultural elements.
@@ Appendix D Retrieval Methodology Details
Appendix E Comparison of K-Viscuit with CVQA dataset
Recent advances in vision-language models have sparked a growing interest in evaluating their cultural awareness across diverse global contexts. A notable contribution in this direction is CVQA (Romero et al., 2024), which introduces a comprehensive multilingual VQA benchmark spanning 28 countries with 9,044 total examples, averaging approximately 323 examples per country. Our K-Viscuit dataset, while focused solely on Korean culture, contains 657 examples—substantially exceeding CVQA’s per-country average. Moreover, CVQA’s Korean subset comprises 290 examples, less than half the size of K-Viscuit. This focused approach enables K-Viscuit to provide more comprehensive coverage of Korean cultural elements, offering deeper insights into VLMs’ understanding of specific cultural contexts.